Solve the system by addition or substitution.
–4x + 2y = –12
y = 2x – 7

Answer by funmath(2933)

Solve the system by addition or substitution.
–4x + 2y = –12
y = 2x – 7
Because one of the equations is solved for a variable already, substitution will lead to less careless errors.
Substitute y=2x-7 into the first equation fo y and solve for x.
-4x+2(2x-7)=-12
-4x+4x-14=-12
-14=-12
Because your variable was eliminated and -14 does no = -12, this system is inconsistent and has no solution. Graphically these two line would be parallel and never intersect.
Happy Calculating!!!